# A Bregman ADMM for Bethe variational problem

### v2

@@ 3.1 Algorithmic framework
In this subsection, we present our algorithm for solving (BVPτ). The alternating direction method of multipliers (ADMM) [20, 22] is widely used in optimization due to its simplicity and broad applications [11, 37, 70, 68]. The traditional ADMM algorithm for solving (BVPτ) is outlined in Algorithm 1.
@@ 3. $\Lambda^{t+1}:=\Lambda^{t}-\rho(\mathcal{A}({\bf q}^{t+1})+\mathcal{B}({\bf Q}^{t+1}))$
While straightforward, Algorithm 1 uses quadratic penalty terms, which can be problematic because $\mathcal{F}({\bf q})$ is strongly concave with its Hessian’s smallest eigenvalue having the order of $-\Omega(1/\tau)$ in $\mathcal{X}_{\tau}$. When $\tau$ is small, $\rho$ must be chosen very large, $\rho=\Omega(1/\tau)$, to ensure that the subproblem in Step 1 remains convex, which makes the algorithm computationally inefficient. When the subproblem is non-convex, finding its global minimizer can be challenging. Even if specialized algorithms could succeed in identifying the global minimizer, the solution may lie on the boundary, which is not a critical point of (BVPτ).
In [75], Wang and Banerjee proposed a Bregman ADMM, which uses the Bregman divergences $B_{\phi}(\cdot,\cdot)$ and $B_{\psi}(\cdot,\cdot)$ as penalty function and proximal term, as shown in Algorithm 2.
Algorithm 2 Bregman ADMM with linear dual update for (BVPτ) [75]
@@
In Algorithm 2, the subproblem in Step 1 is convex as long as $\rho>0$, contrasting with the stricter requirements in Algorithm 1. The convergence of Bregman-type methods such as the Bregman proximal gradient method [23], stochastic Bregman proximal gradient method [18] and Bregman proximal point method [81] has been well-studied in the literature. However, the theoretical analysis of Bregman ADMM is still in its infancy. Although Wang and Banerjee proved the convergence of their Bregman ADMM for convex problems, its extension to non-convex problems like (BVPτ) lacks a theoretical guarantee. While nonconvex ADMM variants have been studied [74, 61, 44], they often use quadratic penalties, with the Bregman divergence appearing only as a proximal term. Here, we refer to a Bregman ADMM as an ADMM that uses Bregman divergence as the penalty function, and the convergence of such a method for non-convex problems remains an open question.
@@
  • The asymmetry of the Bregman divergence penalty function complicates the convergence proofs. In contrast, the convergence proofs of the usual ADMM that employs a quadratic penalty function do not encounter such difficulties. This fact alone hinders a direct adaptation of the proof in Wang and Banerjee [75] for the convergence of convex Bregman ADMM.
  • The marginal $\mathcal{G}({\bf Q})$ defined in (14) is not Lipschitz-continuous on the boundary of $\mathcal{Y}$ defined in (12). However, in existing analyses of non-convex ADMM [76, 80, 25, 45], the objective function (whose variable is optimized) in the second subproblem is typically assumed to have a Lipschitz-continuous gradient (When there are more than two blocks, this assumption is applied to the last function.), which ensures that the Lagrangian multiplier can be bounded in terms of the primal variables.
@@ 3.2 Lemmas for the convergence analysis of Algorithm 3
Because ${\bf Q}^{t}\in\operatorname{ri}(\mathcal{Y}),$ $\mathcal{B}({\bf Q}^{t})\in\mathcal{B}(\operatorname{ri}(\mathcal{Y})).$ In addition, let $X\in\mathcal{Y}$ be such that $X_{ij}:=q_{i}^{t}{q_{j}^{t}}^{\top}.$ From ${\bf q}^{t}\in\operatorname{ri}(\mathcal{X})$ and (13), we have that $X\in\operatorname{ri}(\mathcal{Y})$ and $\mathcal{A}({\bf q}^{t})+\mathcal{B}(X)=0.$ This implies that $-\mathcal{A}({\bf q}^{t})\in\mathcal{B}(\operatorname{ri}(\mathcal{Y})).$ We summarize the results as follows:
@@
Because $\phi(x)+\delta_{[0,1]}(x)=\left\langle x,\log x\right\rangle+\delta_{[0,1]}(x)$ is strongly convex with parameter $1,$ $\hat{\phi}$ is also strongly convex with parameter 1. For any $x\in\mathcal{B}(\operatorname{ri}(\mathcal{Y})),$ because $0<x<1,$ $\hat{\phi}(x)$ is differentiable with gradient $\hat{\mathcal{P}}\left(\nabla\phi(x)\right).$ From the property of a strongly convex function in Proposition 12.6 of [65], we have that for any $x,y\in\mathcal{B}(\operatorname{ri}(\mathcal{Y})),$
@@
$-|\mathcal{E}|M\leq\left\langle{\bf C},{\bf Q}\right\rangle\leq|\mathcal{E}|M,\ -2|\mathcal{E}|\log r\leq\sum_{ij\in\mathcal{E}}\left\langle Q_{ij},\log Q_{ij}\right\rangle\leq 0$
@@
Substituting (3.2) into (68), and using (67) and the property $B_{\phi}(x,y)\geq\frac{1}{2}\|x-y\|^{2}$ for $x,y$ in the relative interior of the unit simplex [7, Prop. 5.1], we get
$-\|\hat{\mathcal{P}}(\Lambda^{t})\|\cdot\|\mathcal{B}({\bf Q}^{t+1})+\mathcal{A}({\bf q}^{t+1})\|+\frac{\rho}{2}\|\mathcal{A}({\bf q}^{t+1})+\mathcal{B}({\bf Q}^{t+1})\|^{2}\leq 2|\mathcal{E}|\cdot\left(M+\log r\right).$ (70)
Now, we use mathematical induction to prove ($i$), ($ii$), ($iii$) and ($iv$). When $t^{\prime}=1,$ $(i)$ holds because $\Lambda^{1}={\bf 0}.$ Suppose $(i)$ holds for $t^{\prime}=t\in\mathbb{N}^{+}.$ We will proceed with the following 4 steps.
@@
$-\xi_{r,\tau}\cdot\|\mathcal{B}({\bf Q}^{t+1})+\mathcal{A}({\bf q}^{t+1})\|+\frac{\rho}{2}\|\mathcal{A}({\bf q}^{t+1})+\mathcal{B}({\bf Q}^{t+1})\|^{2}\leq 2|\mathcal{E}|\cdot\left(M+\log r\right),$ (71)
@@
where the last inequality comes from (65). (3.2) implies that ($iii$) holds for $t+1$.
@@
$[\nabla\phi(x)](i)-[\nabla\phi(y)](i)=\log x(i)-\log y(i)=\int_{y(i)}^{x(i)}\frac{{\rm d}t}{t}.$ (83)
@@
where “././. /” is element-wise division. Similarly, we have that
@@
[…]
@@
where the first inequality comes from the $\frac{2}{\tau}-$Lipschitz continuity of $\log(\cdot)$ in $[\frac{\tau}{2},\infty).$ Substituting (3.3) into (3.3), we get for any $t\geq 2,$
@@
where “././. /” means element-wise division. From (3.3), we have that
@@
where $\xi_{r,\tau}$ is defined in (24). We know that $\mathcal{U}$ is compact and every limit point of $({\bf q}^{t},{\bf Q}^{t},\hat{\mathcal{P}}(\Lambda^{t}))$ is inside $\mathcal{U}.$ Because $\Phi$ is constructed from linear functions, logarithm functions and $\mathcal{X}_{\tau},$ $\mathcal{Y}$ are convex polyhedra, we know that $\Phi$ is a tame function [79, 72] that satisfies the KL property mentioned in Subsection 1.7 (see [4, Section 4]). Because $\mathcal{U}$ is compact, from Lemma 6 in [10], $\Phi$ satisfies the uniformed KL property, that is, there exists $\epsilon,\eta>0$ and a continuous concave function $\varphi:[0,\eta)\rightarrow\mathbb{R}_{+}$ such that for all $\left(\bar{{\bf q}},\bar{{\bf Q}},\bar{\Lambda}\right)\in\mathcal{U}$ with $\Phi(\bar{{\bf q}},\bar{{\bf Q}},\bar{\Lambda})=\Phi^{*}$ and all $({\bf q},{\bf Q},\Lambda)$ in the following set:
@@ 4.1 Solving the subproblem in Step 1
[…]
@@
[…]
@@
Problem (140) is a strongly convex optimization problem with simplex-type constraints, and it is independent for each $k\in[n]$. One can apply the semismooth Newton method [64] to effectively solve its dual problem iteratively.
@@
[…]
@@ 4.2 Solving the subproblem in Step 2
Although Newton or gradient descent methods ensure convergence, they are iterative, resulting in a double-loop structure for the overall algorithm. To create a single-loop variant of Algorithm 3, we add a proximal term to linearize the penalty function in (144). Using the strong subadditivity (SSA) of the entropy function [40], we define the following convex function $\varphi:\mathcal{Y}\to\mathbb{R}$ such that for any ${\bf Q}\in\mathcal{Y},$
@@ 4.4 Penalty parameter
In our theoretical analysis of Algorithm 3, we fix the penalty parameter $\rho$ to be some sufficiently large number. However, in practice, the efficiency of ADMM is sensitive to the penalty parameter and a commonly used strategy is to adaptively update it to balance the primal and dual KKT residuals [43, 70]. In our implementation, we update $\rho^{t}>0$ as follows:
@@ 5.1 Quantum Bethe variational problem
In (5.1), $\widehat{C}_{ij}\in\mathbb{H}^{r^{2}},$ $\hat{c}_{k}\in\mathbb{H}^{r}$ are hermitian coefficient matrices. The operator ${\rm logm}(\cdot)$ refers to the logarithm of a matrix such that for any matrix $A\in\mathbb{H}^{p}_{+}$ (the set of $p\times p$ hermitian positive semidefinite matrix.) with eigenvalue decomposition $P{\rm diag}(d)P^{\top},$ ${\rm logm}(A)=P{\rm diag}(\log(d))P^{\top}.$ Note that ${\rm logm}(A)$ requires $A$ to be positive definite. However, by defining $\left\langle A,{\rm logm}(A)\right\rangle:=\left\langle d,\log(d)\right\rangle,$ $A$ can have zero eigenvalues. In this case, the domain of (5.1) is a compact set. Similarly, we use ${\rm expm}(\cdot)$ to denote the matrix exponential such that ${\rm expm}(A)=P{\rm diag}(\exp(d))P^{\top}.$ The mappings $\mathrm{Tr_{l}}:\mathbb{H}^{r^{2}}\rightarrow\mathbb{H}^{r}$, $\mathrm{Tr_{r}}:\mathbb{H}^{r^{2}}\rightarrow\mathbb{H}^{r}$ are trace out operators such that for any $B\in\mathbb{H}^{r^{2}}$ and $i,j\in[r],$
@@
Here for simplicity, we only provide the optimization formulation of (5.1). For more bankground knowledge about quantum mechanics, please refer to [35, 88]. (5.1) is an extension of (BVP) with the following changes

### v3

@@ 3.1 Algorithmic framework
In this subsection, we present our algorithm for solving (BVPτ). The alternating direction method of multipliers (ADMM) [gabay1976dual, glowinski1975approximation] is widely used in optimization due to its simplicity and broad applications [boyd2011distributed, li2016schur, tang2024self, sun2024accelerating]. The traditional ADMM algorithm for solving (BVPτ) is outlined in Algorithm 1.
@@ 3. $\Lambda^{t+1}:=\Lambda^{t}-\rho(\mathcal{A}({\bf q}^{t+1})+\mathcal{B}({\bf Q}^{t+1}))$
While straightforward, Algorithm 1 uses quadratic penalty terms, which can be problematic because $\mathcal{F}({\bf q})$ is strongly concave with its Hessian’s smallest eigenvalue having the order of $\Omega(1/\tau)$ in $\mathcal{X}_{\tau}$. When $\tau$ is small, $\rho$ must be chosen very large, $\rho=\Omega(1/\tau)$, to ensure that the subproblem in Step 1 remains convex, which makes the algorithm computationally inefficient. When the subproblem is non-convex, finding its global minimizer can be challenging. Even if specialized algorithms could succeed in identifying the global minimizer, the solution may lie on the boundary, which is not a critical point of (BVPτ).
In [wang2014bregman], Wang and Banerjee proposed a Bregman ADMM, which uses the Bregman divergences $B_{\phi}(\cdot,\cdot)$ and $B_{\psi}(\cdot,\cdot)$ as penalty function and proximal term, as shown in Algorithm 2.
Algorithm 2 Bregman ADMM with linear dual update for (BVPτ) [wang2014bregman]
@@
In Algorithm 2, the subproblem in Step 1 is convex as long as $\rho>0$, contrasting with the stricter requirements in Algorithm 1. The convergence of Bregman-type methods such as the Bregman proximal gradient method [hanzely2021accelerated], stochastic Bregman proximal gradient method [ding2023nonconvex] and Bregman proximal point method [yang2022bregman] has been well-studied in the literature. However, the theoretical analysis of Bregman ADMM is still in its infancy. Although Wang and Banerjee proved the convergence of their Bregman ADMM for convex problems, its extension to non-convex problems like (BVPτ) lacks a theoretical guarantee. While nonconvex ADMM variants have been studied [wang2018convergence, pham2024bregman, liu2023bregman], they often use quadratic penalties, with the Bregman divergence appearing only as a proximal term. Here, we refer to a Bregman ADMM as an ADMM that uses Bregman divergence as the penalty function, and the convergence of such a method for non-convex problems remains an open question.
@@
  • The asymmetry of the Bregman divergence penalty function complicates the convergence proofs. In contrast, the convergence proofs of the usual ADMM that employs a quadratic penalty function do not encounter such difficulties. This fact alone hinders a direct adaptation of the proof in Wang and Banerjee [wang2014bregman] for the convergence of convex Bregman ADMM.
  • The marginal $\mathcal{G}({\bf Q})$ defined in (14) is not Lipschitz-continuous on the boundary of $\mathcal{Y}$ defined in (12). However, in existing analyses of non-convex ADMM [wang2019global, yang2017alternating, hong2016convergence, liu2019linearized], the objective function (whose variable is optimized) in the second subproblem is typically assumed to have a Lipschitz-continuous gradient (When there are more than two blocks, this assumption is applied to the last function.), which ensures that the Lagrangian multiplier can be bounded in terms of the primal variables.
@@ 3.2 Lemmas for the convergence analysis of Algorithm 3
Because ${\bf Q}^{t}\in\operatorname{ri}(\mathcal{Y}),$ $\mathcal{B}({\bf Q}^{t})\in\mathcal{B}(\operatorname{ri}(\mathcal{Y})).$ In addition, let $X\in\mathcal{Y}$ be such that $X_{ij}:=q_{i}^{t}{q_{j}^{t}}^{\top}.$ From ${\bf q}^{t}\in\operatorname{ri}(\mathcal{X})$ and (13), we have that $X\in\operatorname{ri}(\mathcal{Y})$ and $\mathcal{A}({\bf q}^{t})+\mathcal{B}(X)=0.$ This implies that $\mathcal{A}({\bf q}^{t})\in\mathcal{B}(\operatorname{ri}(\mathcal{Y})).$ We summarize the results as follows:
@@
Because $\phi(x)+\delta_{[0,1]}(x)=\left\langle x,\log x\right\rangle+\delta_{[0,1]}(x)$ is strongly convex with parameter $1,$ $\hat{\phi}$ is also strongly convex with parameter 1. For any $x\in\mathcal{B}(\operatorname{ri}(\mathcal{Y})),$ because $0<x<1,$ $\hat{\phi}(x)$ is differentiable with gradient $\hat{\mathcal{P}}\left(\nabla\phi(x)\right).$ From the property of a strongly convex function in Proposition 12.6 of [rockafellar2009variational], we have that for any $x,y\in\mathcal{B}(\operatorname{ri}(\mathcal{Y})),$
@@
$|\mathcal{E}|M\leq\left\langle{\bf C},{\bf Q}\right\rangle\leq|\mathcal{E}|M,\ -2|\mathcal{E}|\log r\leq\sum_{ij\in\mathcal{E}}\left\langle Q_{ij},\log Q_{ij}\right\rangle\leq 0$
@@
Substituting (3.2) into (68), and using (67) and the property $B_{\phi}(x,y)\geq\frac{1}{2}\|x-y\|^{2}$ for $x,y$ in the relative interior of the unit simplex [beck2003mirror, Prop. 5.1], we get
$\|\hat{\mathcal{P}}(\Lambda^{t})\|\cdot\|\mathcal{B}({\bf Q}^{t+1})+\mathcal{A}({\bf q}^{t+1})\|+\frac{\rho}{2}\|\mathcal{A}({\bf q}^{t+1})+\mathcal{B}({\bf Q}^{t+1})\|^{2}\leq 2|\mathcal{E}|\cdot\left(M+\log r\right).$ (70)
Now, we use mathematical induction to prove ($i$), (iiii), (iiiiii) and ($iv$). When $t^{\prime}=1,$ $(i)$ holds because $\Lambda^{1}={\bf 0}.$ Suppose $(i)$ holds for $t^{\prime}=t\in\mathbb{N}^{+}.$ We will proceed with the following 4 steps.
@@
$\xi_{r,\tau}\cdot\|\mathcal{B}({\bf Q}^{t+1})+\mathcal{A}({\bf q}^{t+1})\|+\frac{\rho}{2}\|\mathcal{A}({\bf q}^{t+1})+\mathcal{B}({\bf Q}^{t+1})\|^{2}\leq 2|\mathcal{E}|\cdot\left(M+\log r\right),$ (71)
@@
where the last inequality comes from (65). (3.2) implies that (iiiiii) holds for $t+1$.
@@
$\nabla\phi(x)](i)-[\nabla\phi(y)](i)=\log x(i)-\log y(i)=\int_{y(i)}^{x(i)}\frac{{\rm d}t}{t}.$ (83)
@@
where “././” is element-wise division. Similarly, we have that
@@
$\rho B_{\phi}\left(\mathcal{B}({\bf Q}^{t+1}),\mathcal{B}({\bf Q}^{t})\right)-\frac{4\sqrt{\rho}\tilde{\xi}_{r,\tau}\|\mathcal{B}({\bf Q}^{t+1})-\mathcal{B}({\bf Q}^{t})\|^{2}}{\tau^{2}}\;\geq\;0,$ (105)
@@
where the first inequality comes from the $\frac{2}{\tau}$-Lipschitz continuity of $\log(\cdot)$ in $\frac{\tau}{2},\infty).$ Substituting (3.3) into (3.3), we get for any $t\geq 2,$
@@
where “././” means element-wise division. From (3.3), we have that
@@
where $\xi_{r,\tau}$ is defined in (24). We know that $\mathcal{U}$ is compact and every limit point of $({\bf q}^{t},{\bf Q}^{t},\hat{\mathcal{P}}(\Lambda^{t}))$ is inside $\mathcal{U}.$ Because $\Phi$ is constructed from linear functions, logarithm functions and $\mathcal{X}_{\tau},$ $\mathcal{Y}$ are convex polyhedra, we know that $\Phi$ is a tame function [wilkie1996model, van1998tame] that satisfies the KL property mentioned in Subsection 1.7 (see [attouch2010proximal, Section 4]). Because $\mathcal{U}$ is compact, from Lemma 6 in [bolte2014proximal], $\Phi$ satisfies the uniformed KL property, that is, there exists $\epsilon,\eta>0$ and a continuous concave function $\varphi:[0,\eta)\rightarrow\mathbb{R}_{+}$ such that for all $\left(\bar{{\bf q}},\bar{{\bf Q}},\bar{\Lambda}\right)\in\mathcal{U}$ with $\Phi(\bar{{\bf q}},\bar{{\bf Q}},\bar{\Lambda})=\Phi^{*}$ and all $({\bf q},{\bf Q},\Lambda)$ in the following set:
@@ 4.1 Solving the subproblem in Step 1
$+\rho\sum_{kj\in\mathcal{E}}\left\langle\log q_{k}-\log(Q_{kj}^{t}{\bf 1}_{r}),q_{k}\right\rangle+\rho\sum_{ik\in\mathcal{E}}\left\langle\log q_{k}-\log({Q_{ik}^{t}}^{\top}{\bf 1}_{r}),q_{k}\right\rangle:\ {\bf 1}^{\top}_{r}q_{k}=1,\ q_{k}\geq\tau\Big\}.$
@@
$q^{t+1}_{k}:=\arg\min\Big\{\left\langle\hat{c}_{k}^{t},q_{k}\right\rangle+\rho\cdot d_{k}\left\langle q_{k},\log q_{k}\right\rangle:\ {\bf 1}^{\top}_{r}q_{k}=1,\ q_{k}\geq\tau\Big\},$ (140)
@@
Problem (140) is a strongly convex optimization problem with simplex-type constraints, and it is independent for each $k\in[n]$. One can apply the semismooth Newton method [qi2010semismooth] to effectively solve its dual problem iteratively.
@@
$q^{t+1}_{k}:=\arg\min\Big\{\left\langle\hat{c}_{k}^{t},q_{k}\right\rangle+\rho\cdot d_{k}\left\langle q_{k},\log q_{k}\right\rangle:\ {\bf 1}^{\top}_{r}q_{k}=1,\ q_{k}\geq 0\Big\},$ (142)
@@ 4.2 Solving the subproblem in Step 2
Although Newton or gradient descent methods ensure convergence, they are iterative, resulting in a double-loop structure for the overall algorithm. To create a single-loop variant of Algorithm 3, we add a proximal term to linearize the penalty function in (144). Using the strong subadditivity (SSA) of the entropy function [lieb2002some], we define the following convex function $\varphi:\mathcal{Y}\to\mathbb{R}$ such that for any ${\bf Q}\in\mathcal{Y},$
@@ 4.4 Penalty parameter
In our theoretical analysis of Algorithm 3, we fix the penalty parameter $\rho$ to be some sufficiently large number. However, in practice, the efficiency of ADMM is sensitive to the penalty parameter and a commonly used strategy is to adaptively update it to balance the primal and dual KKT residuals [lin2011linearized, tang2024self]. In our implementation, we update $\rho^{t}>0$ as follows:
@@ 5.1 Quantum Bethe variational problem
In (5.1), $\widehat{C}_{ij}\in\mathbb{H}^{r^{2}},$ $\hat{c}_{k}\in\mathbb{H}^{r}$ are hermitian coefficient matrices. The operator ${\rm logm}(\cdot)$ refers to the logarithm of a matrix such that for any matrix $A\in\mathbb{H}^{p}_{+}$ (the set of $p\times p$ hermitian positive semidefinite matrix.) with eigenvalue decomposition $P{\rm diag}(d)P^{\top},$ ${\rm logm}(A)=P{\rm diag}(\log(d))P^{\top}.$ Note that ${\rm logm}(A)$ requires $A$ to be positive definite. However, by defining $\left\langle A,{\rm logm}(A)\right\rangle:=\left\langle d,\log(d)\right\rangle,$ AA can have zero eigenvalues. In this case, the domain of (5.1) is a compact set. Similarly, we use ${\rm expm}(\cdot)$ to denote the matrix exponential such that ${\rm expm}(A)=P{\rm diag}(\exp(d))P^{\top}.$ The mappings $\mathrm{Tr_{l}}:\mathbb{H}^{r^{2}}\rightarrow\mathbb{H}^{r}$, $\mathrm{Tr_{r}}:\mathbb{H}^{r^{2}}\rightarrow\mathbb{H}^{r}$ are trace out operators such that for any $B\in\mathbb{H}^{r^{2}}$ and $i,j\in[r],$
@@
Here for simplicity, we only provide the optimization formulation of (5.1). For more bankground knowledge about quantum mechanics, please refer to [leifer2008quantum, zhao2024quantum]. (5.1) is an extension of (BVP) with the following changes